David​ Ricardo, the British political​ economist, used the example of two commodities —wine

and cloth —produced by England and Portugal to explain trade. The following table shows the number of labor hours it would take England and Portugal to produce one unit each of wine and​ cloth:

Portugal

England

Wine

120

80

Cloth

100

90

Portugal can produce both wine and cloth using fewer labor hours than England. A group of​ Mercantilists, who believe that nations build their wealth by exporting more than they​ import, suggest that Portugal has nothing to gain from trading with England.

Would you​ agree?

A.

​Yes, because if they both produce the same two​ goods, there's no point in trading with England.

B.

​No, gains from trade are determined on the basis of comparative​ advantage, not absolute advantage.

C.

​Yes, Portugal produces both wine and cloth using fewer labor hours than​ England, so it has an absolute advantage in the production of both goods and therefore can produce more efficiently for itself.

D.

​No, because to export they have to​ trade, and so they must trade with England.

Suppose that each country spends 100 hours producing each good. Fill in the table below to show how much of each good would be produced.

​(Round

your responses to two decimal

places.​)

Portugal

England

Wine

____

​unit(s)

____

​unit(s)

Cloth

____

​unit(s)

____

​unit(s)
Answer #1

1) Option B.

​No, gains from trade are determined on the basis of comparative​ advantage, not absolute advantage.

2) With 100 hours each country can produce

Portugal

England

Wine

100/120=

​0.83 unit(s)

100/80

​1.25 unit(s)

Cloth

100/100=

​1 unit(s)

100/90=

​1.11 unit(s)